# Generalized local jocabians and commutative group stacks

### v2

@@ Abstract.
In [CS01, Page 109] Grothendieck sketches the construction of a complex $J_{*}(X)$ made of commutative pro-algebraic groups, associated to a smooth variety $X$, and for which each $J_{i}(X)$ is a product of local factors called the local generalized jacobians. The complex $J_{*}(X)$ has the property to compute Zariski cohomology of $X$ with coefficients in algebraic groups and play the role of a homology of $X$. The purpose of this note is to recast this construction in the setting of higher algebraic group stacks for the fppf topology instead. For this, we introduce a notion of algebraic homology associated to a scheme which is a universal object computing fppf cohomology with coefficients in group schemes. We endow this algebraic homology with a filtration by dimension of supports, and prove that, when $X$ is smooth, $J_{*}(X)$ appears as the $E_{1}$-page of the corresponding spectral sequence. In a final part we partially extends our constructions and results over arbitrary bases.
@@ Introduction
The main purpose of this note is to introduce a notion of algebraic homology, associated to schemes over a perfect field $k$, which is a universal object computing flat cohomology with coefficients in various commutative group schemes. The idea of algebraic homology is certainly not new, and was already sketched in a slight different setting in a letter of Grothendieck to Serre (see letter from 8.9.1960 in [CS01, page 109]), as well as in [Gro] (cote 134-2 page 60) and in [Mil76] for instance. However, in the present work we realize algebraic homology as a universal (pro)object in the $\infty$-category of commutative groups in algebraic higher stacks, whereas in the above references it is constructed as a complex of commutative pro-algebraic groups. The main result of this work is a comparison statement for smooth schemes, and states that our notion of algebraic homology recovers Grothendieck’s orginal construction. This comparison is not a formal statement as it relates two objects living in two non-equivalent categories, and requires to use an extra piece of data in the form of a filtration by dimension of supports. Grothendieck’s original construction is shown to appear as the 00-th truncation of algebraic homology, for the Beilinson’s t-structure on filtered objects (see for instance [Pav23]). This is a fancy manner to states that Grothendieck’s complex of pro-algebraic groups appears as the $E_{1}$-page of the spectral sequence associated to the filtration by dimension of supports on algebraic homology. We also show that both objects can be used in order to compute fppf cohomology with coefficients in commutative group schemes (with some affineness and unipotency conditions for higher cohomology groups). This last statement is true by definition for algebraic homology but requires a proof in the case of Grothendieck’s construction which is done by gluing together the so-called local generalized jacobians appearing in our title.
We work over a perfect field $k$, and in the $\infty$-topos of hypercomplete fppf stacks $\mathbf{St}_{k}$. In a first part of this work we introduce the notions of commutative algebraic group stacks over $k$ (see definition 1.3), which are the commutative group objects inside the $\infty$-category of Artin stacks locally of finite type over $k$ and whose diagonal are quasi-compact (in a strong sense). These are higher stacks analogues of locally algebraic groups over $k$, that is separated commutative group schemes locally of finite type. The commutative algebraic group stacks can also be characterized as connective complexes of sheaves of abelian groups $E$ of the big fppf site, whose cohomology sheaves $H^{i}(E)$ are all representable by locally algebraic groups, and are moreover quasi-projective if $i<0$. The algebraic homology of a scheme $X$ over $k$ (or any stack over $k$) is then defined as follows (see definition 3.1).
@@
where $F_{i}\mathbb{H}^{alg}(X)$ is the part of the algebraic homology which is supported in codimension $i$ in $X$. The $d$-th graded piece of this filtration is given by the product of all local algebraic homologies over all the points of codimension $d$ in $X$ (see proposition 3.2)
$Gr_{d}\mathbb{H}^{alg}(X)\simeq\prod_{x\in X^{(dimX-d)}}\mathbb{H}^{alg}_{x}(X).$
When $X$ is furthermore smooth over $k$, purity implies that each graded piece $Gr_{d}\mathbb{H}^{alg}(X)$ is in fact $(n-d)$-connective (see corollary 3.4). In another word, the filtered object $F_{*}\mathbb{H}^{alg}(X)$ is 00-connective with respect to Beilinson’s t-structure on filtered objects. As being connective, we get a canonical projection of filtered objects
@@
where $H^{0}_{B}$ refers to the 00-th cohomology object for the Beilinson t-structure. The object $H^{0}_{B}(F_{*}\mathbb{H}^{alg}(X))$ is a genuine complex made of commutative pro-algebraic groups, and as such can be compared to Grothendieck’s original construction $J_{*}(X)$. In the complex $J_{*}(X)$, the degree $d$ part is the product of local generalized jacobian $J_{x}(X)$ over all points $x$ of codimension $d$ in $X$ which are defined to pro-represent local cohomology in degree $d$ (see definition 3.5).
Our main comparison result is the following theorem.
@@
We have a natural isomorphism of complexes of pro-algebraic groups over $k$
@@
The theorem below, shows that $J_{*}(X)$ appears as the 00-th truncation of $F_{*}\mathbb{H}^{alg}(X)$. However, the natural projection $F_{*}\mathbb{H}^{alg}(X)\to J_{*}(X)$ is not an equivalence in general, and it is thus surprising that both objects still compute the same cohomologies with unipotent coefficients.
@@
$Ext^{i}_{naive}(J_{*}(X),G)\simeq Ext^{i}_{fppf}(\mathbb{H}^{alg}(X),G)\simeq H^{i}_{fppf}(X,G).$
@@
In the theorem above the $Ext^{i}_{naive}$ refers to the Yoneda ext-groups computed inside the derived category of complexes of commutative pro-algebraic groups. These ext-groups differ from the ext-groups computed as fppf sheaves, and thus are not given by ext-groups computed inside our category of commutative group stacks. We note that the theorem above is already stated in [CS01, page 109] where the fppf topology is replaced with the Zariski topology.
@@
All the results mentioned above are true over a perfect field $k$. We do not know a way to extend our results and constructions over more general base schemes, at least in an interesting manner. The reader will notice that a key ingredient for our results is purity for local cohomology of regular schemes with coefficients in group schemes. Very general purity theorems have been proven recently in [CS24], but we would rather need purity results with coefficients in commutative algebraic group stacks, and we do not see a simple manner to obtained those from the case of flat group schemes. We hope to be able to come back to this question in a future work.
@@
However, when restricted to unipotent coefficients, we propose a generalization of algebraic homology over any base ring, based on our notion of affine stacks (see [Toe06]) and their abelian counter-parts which we call affine homology types. We prove that any (small) stack $X$, in particular any scheme, maps to a universal affine homology types called its affine homology and denoted by $\mathbb{H}^{u}(X)$ (see proposition 4.6 and definition 4.8). Over a field, $\mathbb{H}^{u}(X)$ is the unipotent completion of algebraic homology, and its advantage is that it makes sense over any base. Affine homology comes equipped with a filtration by dimension of supports whose associated graded are given by local affine homologies. Moreover, we prove that $\mathbb{H}^{u}(X)$ is completely characterized by the complex $\mathbb{R}\Gamma(X,\mathbb{G}_{a})$ considered as a dg-algebra over $\mathbb{R}End(\mathbb{G}_{a})$, which we consider as a form of Dieudonné theory for affine homology types.
@@
To finish this introduction we would like to mention several related works. As we already mentioned the notion of algebraic homology already appears in [CS01, page 109]) and [Mil76]. The results of the present work does not pretend to be very original and the main point of this note was rather to recast these works in the more modern language of higher stacks. The only exception is possibly our last section on affine homology, which we think can be of independent interest. After posting a first version of this work O. Gwilliam pointed to us the work [Mir17] in which a general setting for algebraic homology based on higher commutative group stacks is presented.
@@
In another direction, the results of this work have some limitations. First of all the fact that we work over a base field is restrictive. Over more general bases, the $1$-truncation of algebraic homology essentially appears in the work of S. Brochard in [Bro21]. Extending the results of [Bro21] seems however difficult because of the presence of complicated higher ext groups between commutative group schemes which prevent to a perfect duality by mapping to $\mathbb{G}_{m}$. Also, the approach in [Bro21] is based on duality, which is an aspect we havent touched at all in this note. Grothendieck already mentions duality in [Gro, 134-2], and in degree 00 this duality of local generalized jacobians have been studied in [CC90] (it also appears in [Mir17]).
@@
Finally, there are also relations with algebraic homotopy types, as studies for instance in [Toe06, KPT09, BV15, MR23]. Our notion of affine homology is the ablelian analogue of these, and the two should be related by means of the Hurewitz morphism. As a final comment, affine stacks are duals to cosimplicial commutative algebras, also called $LSym$-algebras. These are Koszul dual to partition Lie algebras, thanks to [BM23], and therefore our notion of affine homology should be closely related to that of abelian partition Lie algebras.
Acknowledgments: In a first version of this work algebraic homology was only studied with some extra artificial unipotency conditions. I warmly thank B. Kahn for his comments during a lecture on the content of this work which has convinced me that these extra conditions were not necessary at all.
@@ 1. Commutative algebraic group stacks
We work over a perfect base field $k$. We denote by $\mathbf{Aff}_{k}$ the category of affine $k$-schemes, which will be endowed with the fppf topology. The $\infty$-category of (hypercomplete) stacks on $\mathbf{Aff}_{k}$ is denoted by $\mathbf{St}_{k}$.
@@ 1.1. Reminders on abelian group schemes
We denote by $\mathbf{AbSch}_{k}$, the category of commutative group schemes over $k$ which are separated and of finite type over $k$ (or equivalently quasi-projective over $k$, see for instance [sta] lemma 39.8.2, tag 0BF6). The Yoneda embedding provides a fully faithful functor
@@
from $\mathbf{AbSch}_{k}$ to the category $\mathbf{Ab}_{k}$ of sheaves of abelian groups on $\mathbf{Aff}_{k}$ endowed with the $fppf$-topology. It is well known that this full embedding commutes with the formations of kernels and cokernels, and that its essential image is also stable by extensions. We do not know a precise reference for this, so we include it as a lemma.
@@
The full sub-category of $\mathbf{Ab}_{k}$ formed by sheaves which are representable by commutative separated group schemes of finite type over $k$, is stable by taking kernels, cokernels and extensions.
Proof. The stability by kernels is obvious as the Yoneda embedding is left exact. For the stability by cokernels, the quotient fppf-sheaf $G/H$, for a monomorphism of commutative algebraic groups $H\hookrightarrow G$, is always representable by an Artin stack whose diagonal is a monomorphism, that is by an algebraic space. This algebraic space is moreover quasi-separated (because $H$ is quasi-compact) and of finite type. We can thus use Artin’s theorem stating that a group object in the category of quasi-separated algebraic spaces over a field is representable by a (separated) group scheme (see [Art69, Lem. 4.2]). Finally, if a sheaf of abelian groups is an extension of algebraic groups, then it is, by descent, representable by a quasi-separated algebraic space, and we can conclude using again Artin’s result. $\Box$
The Barsotti-Chevalley decomposition theorem (see [Mil13]) states that any object $G\in\mathbf{AbSch}_{k}$ fits in an exact sequence
@@
where $G_{0}$ is affine and $A$ is an abelian variety. Moreover, by [DG70, II-5 Cor. 2.3] $G_{0}$ itself sits in an exact sequence $0\to K\to G_{0}\to H\to 0,$ where $K$ is smooth and affine, and $H$ is finite infinitesimal (i.e. $H_{red}=0$). We also recall from [DG70, IV-2 Porp. 2.5] that $G\in\mathbf{AbSch}_{k}$ is unipotent if it admits a finite filtration by sub-group schemes
@@
such that each $G_{i}/G_{i+1}$ can be realized as a sub-group scheme the additive group $\mathbb{G}_{a}$. We can moreover arrange things so that each $G_{i}/G_{i+1}$ is a twisted form of $\mathbb{Z}/p$ or of $\alpha_{p}$. Unipotent group schemes are automatically affine, and form a full sub-category of $\mathbf{AbSch}_{k}$ stable by taking kernels, cokernels and extensions.
@@ 1.2. Reminders on purity for flat cohomology
We remind here the purity theorem we will use in a crucial manner in order to compare algebraic homology with Grothendieck’s original definition. The most general results on purity can be found in [CS24], but for smooth schemes $X$ over a perfect $k$ these can be proven with simple technology, which we recall below.
@@
Case $G=\mathbb{G}_{m}$. For $d=n$ there is nothing to prove, and when $d=n-1$ the statement simply follows from the fact that $X_{x}-\{x\}$ is dense in $X_{x}$. We thus assume that $d<n-1$. The case $i=1$ is true by Hartogs, as the restriction map $Pic(X_{x})\to Pic(X_{x}-\{x\})$ is an equivalence of groupoids. For $i=2$ the proposition follows from the fact that $H^{2}_{et}(X_{x},\mathbb{G}_{m})\to H^{2}_{et}(X_{x}-\{x\},\mathbb{G}_{m})$ is always injective (see [Gro68, Cor. 1.10]). When $i>2$, the fact that $\mathbb{G}_{m}$ has torsion cohomology in degree $2$ or higher (see [Gro68, Prop. 1.4]) implies that $H^{i}_{x}(X,\mathbb{G}_{m})$ is always torsion. We are thus led to prove the statement for the group $\mu_{n}$ for some $n$, and by devissage we can even assume that $n$ is prime. If $n$ is invertible in $k$, then the purity theorem in étale cohomology (see for instance [Mil80, §VI Thm. 5.1]) implies the result. When $n=p$ is the characteristic of the base field, we use the following standard "resolution" of $\mu_{p}$: on the small fppf site of $X$, there is an equivalence
@@
where $\Omega_{X/k}^{1,cl}\subset\Omega_{X/k}^{1}$ is the subsheaf of closed $1$-forms and sits in cohomological degrees 00 (see [Mil80, §III Prop. 4.14]). The differential $\Omega_{X/k}^{1,cl}\to\Omega_{X/k}^{1}$ is here $i-C$, where $i$ is the standard inclusion and $C$ is the composition of the projection $\Omega_{X/k}^{1,cl}\to H^{1}_{DR}(X/k)$ and the Cartier isomorphism $C:H^{1}_{DR}(X/k)\simeq\Omega^{1}_{X/k}$. Both abelian sheaves $\Omega_{X/k}^{1,cl}$ and $\Omega_{X/k}^{1}$ are quasi-coherent (the first one on the Frobenius twists $X^{(1)}$ and the second one on $X$), and thus satisfy purity. We deduce from this that $H^{i}_{x}(X,\mu_{p}[1])\simeq H^{i+1}_{x}(X,\mu_{p})\simeq 0$ as soon as $i<d$ and in particular that $H^{i}_{x}(X,\mu_{p})\simeq 0$ when $i<d$.
@@
1.3. Commutative algebraic group stacks

### v3

# Generalized local jacobians and commutative group stacks

@@ Abstract.
In [10, page 109] Grothendieck sketches the construction of a complex $J_{*}(X)$ made of commutative pro-algebraic groups, associated to a smooth variety $X$, and for which each $J_{i}(X)$ is a product of local factors called the local generalized jacobians. The complex $J_{*}(X)$ has the property to compute Zariski cohomology of $X$ with coefficients in algebraic groups and plays the role of a homology of $X$. The purpose of this note is to recast this construction in the setting of higher algebraic group stacks for the fppf topology instead. For this, we introduce a notion of algebraic homology associated to a scheme over a perfect field which is a universal object computing fppf cohomology with coefficients in group schemes. We endow this algebraic homology with a filtration by dimension of supports, and prove that, when $X$ is smooth of dimension $n$, it is $n$-connective with respect to Beilinson’s t-structure on the category of filtered objects, and that $J_{*}(X)$ appears as its $n$-th truncation. In a final part we partially extend our constructions and results over arbitrary bases.
@@ Introduction
The main purpose of this note is to introduce a notion of algebraic homology, associated to schemes over a perfect field $k$, which is a universal object computing flat cohomology with coefficients in various commutative group schemes. The idea of algebraic homology is certainly not new, and was already sketched in a slightly different setting in a letter of Grothendieck to Serre (see letter from 8.9.1960 in [10, page 109]), as well as in [14] (cote 134-2 page 60) and in [20] for instance. However, in the present work we realize algebraic homology as a universal (pro)object in the $\infty$-category of commutative groups in algebraic higher stacks, whereas in the above references it is constructed as a complex of commutative pro-algebraic groups. The main result of this work is a comparison statement for smooth schemes, and states that our notion of algebraic homology recovers Grothendieck’s original construction. This comparison is not a formal statement as it relates two objects living in two non-equivalent categories, and requires to use an extra piece of data in the form of a filtration by dimension of supports. Grothendieck’s original construction is shown to appear as the $n$-th truncation of algebraic homology, for the Beilinson’s t-structure on filtered objects (see for instance [27]). We also show that both objects can be used in order to compute fppf cohomology with coefficients in commutative group schemes (with some affineness and unipotency conditions for higher cohomology groups). This last statement is true by definition for algebraic homology but requires a proof in the case of Grothendieck’s construction which is done by gluing together the so-called local generalized jacobians appearing in our title.
We work over a perfect field $k$, and in the $\infty$-topos of hypercomplete fppf stacks $\mathbf{St}_{k}$. In a first part of this work we introduce the notions of commutative algebraic group stacks over $k$ (see Definition 1.3), which are the commutative group objects inside the $\infty$-category of Artin stacks locally of finite type over $k$ and whose diagonal is quasi-compact (in a strong sense). These are higher stacks analogues of locally algebraic groups over $k$, that is commutative group schemes locally of finite type. The commutative algebraic group stacks can also be characterized as connective complexes of sheaves of abelian groups $E$ of the big fppf site, whose cohomology sheaves $H^{i}(E)$ are all representable by locally algebraic groups, and are moreover quasi-projective if $i<0$. The algebraic homology of a scheme $X$ over $k$ (or any stack over $k$) is then defined as follows (see Definition 2.1).
@@
where $F_{i}\mathbb{H}^{alg}(X)$ is the part of the algebraic homology which is supported in dimension $i$ in $X$ (see definition 3.1). The $d$-th graded piece of this filtration is given by the product of all local algebraic homologies over all the points of dimension $d$ in $X$ (see Proposition 3.2)
$Gr_{d}\mathbb{H}^{alg}(X)\simeq\prod_{x\in X^{(d)}}\mathbb{H}^{alg}_{x}(X_{x}).$
When $X$ is furthermore smooth over $k$, purity implies that each graded piece $Gr_{d}\mathbb{H}^{alg}(X)$ is in fact (DimXd)(DimX-d)-connective (see Corollary 3.4). In other words, the filtered object $F_{*}\mathbb{H}^{alg}(X)$ is 0-connective with respect to Beilinson’s t-structure on filtered objects. Being connective, we get a canonical projection of filtered objects
@@
where $H^{0}_{B}$ refers to the 0-th cohomology object for the Beilinson t-structure. The object $H^{0}_{B}(F_{*}\mathbb{H}^{alg}(X))$ is a genuine complex made of commutative pro-algebraic groups, and as such can be compared to Grothendieck’s original construction $J_{*}(X)$. In the complex $J_{*}(X)$, the degree $d$ part is the product of local generalized jacobian $J_{x}(X)$ over all points $x$ of codimension $d$ in $X$ which are defined to pro-represent local cohomology in degree $d$ (see Definition 3.5).
Our main comparison result is the following theorem (see Corollary 3.6).
@@
If $X$ is smooth over $k$, we have a natural isomorphism of complexes of pro-algebraic groups over $k$
@@
The theorem below, shows that $J_{*}(X)$ appears as the 0-th truncation of $F_{*}\mathbb{H}^{alg}(X)$. However, the natural projection $F_{*}\mathbb{H}^{alg}(X)\to J_{*}(X)$ is not an equivalence in general, and it is thus surprising that both objects still compute the same cohomologies with unipotent coefficients.
@@
$Ext^{i}_{naive}(J_{*}(X),G)\simeq H^{i}_{fppf}(X,G).$
@@
In the theorem above the $Ext^{i}_{naive}$ refers to the Yoneda ext-groups computed inside the derived category of complexes of commutative pro-algebraic groups. These ext-groups differ from the ext-groups computed as fppf sheaves, and thus are not given by ext-groups computed inside our category of commutative group stacks. We note that the theorem above is already stated in [10, page 109] where the fppf topology is replaced with the Zariski topology.
@@
All the results mentioned above are true over a perfect field $k$. We do not know a way to extend our results and constructions over more general base schemes, at least in an interesting manner. The reader will notice that a key ingredient for our results is purity for local cohomology of regular schemes with coefficients in group schemes. Very general purity theorems have been proven recently in [8], but we would rather need purity results with coefficients in commutative algebraic group stacks, and we do not see a simple manner to obtain those from the case of flat group schemes. We hope to be able to come back to this question in a future work.
@@
However, when restricted to unipotent coefficients, we propose a generalization of algebraic homology over any base ring, based on our notion of affine stacks (see [32]) and their abelian counter-parts which we call affine homology types. We prove that any (small) stack $X$, in particular any scheme, maps to a universal affine homology type called its affine homology and denoted by $\mathbb{H}^{u}(X)$ (see Proposition 4.6 and Definition 4.8). Over a field, $\mathbb{H}^{u}(X)$ is the unipotent completion of algebraic homology, and its advantage is that it makes sense over any base. Affine homology comes equipped with a filtration by dimension of supports whose associated graded pieces are given by local affine homologies. Moreover, we prove that $\mathbb{H}^{u}(X)$ is completely characterized by the complex $\mathbb{R}\Gamma(X,\mathbb{G}_{a})$ considered as a dg-algebra over $\mathbb{R}End(\mathbb{G}_{a})$, which we consider as a form of Dieudonné theory for affine homology types.
@@
To finish this introduction we would like to mention several related works. As we already mentioned the notion of algebraic homology already appears in [10, page 109]) and [20]. The results of the present work do not pretend to be very original and the main point of this note was rather to recast these works in the more modern language of higher stacks. The only exception is possibly our last section on affine homology, which we think can be of independent interest. After posting a first version of this work O. Gwilliam pointed to us the work [23] in which a general setting for algebraic homology based on higher commutative group stacks is presented.
@@
In another direction, the results of this work have some limitations. First of all the fact that we work over a base field is restrictive. Over more general bases, the $1$-truncation of algebraic homology essentially appears in the work of S. Brochard in [7]. Extending the results of [7] seems however difficult because of the presence of complicated higher ext groups between commutative group schemes which prevent to a perfect duality by mapping to $\mathbb{G}_{m}$. Also, the approach in [7] is based on duality, which is an aspect we haven’t touched at all in this note. Grothendieck already mentions duality in [14, 134-2], and in degree 0 this duality of local generalized jacobians has been studied in [11] (it also appears in [23]).
@@
Finally, there are also relations with algebraic homotopy types, as been studied for instance in [32, 17, 2, 24]. Our notion of affine homology is the abelian analogue of these, and the two should be related by means of the Hurewitz morphism. As a final comment, affine stacks are duals to cosimplicial commutative algebras, also called $LSym$-algebras. These are Koszul dual to partition Lie algebras, thanks to [3], and therefore our notion of affine homology should be closely related to that of abelian partition Lie algebras.
Acknowledgments: In a first version of this work algebraic homology was only studied with some extra artificial unipotency conditions. I warmly thank B. Kahn for his comments during a lecture on the content of this work which has convinced me that these extra conditions were not necessary at all. I am also very grateful to the anonymous referee for his very detailed and very useful comments.
@@ 1. Commutative algebraic group stacks
We work over a perfect base field $k$. We denote by $\mathbf{Aff}_{k}$ the category of affine $k$-schemes, which will be endowed with the fppf topology.
@@ 1.1. Reminders on abelian group schemes
We denote by $\mathbf{AbSch}_{k}$, the category of commutative group schemes of finite type over $k$ (or equivalently quasi-projective over $k$, see for instance [28, tag 0BF6]). The Yoneda embedding provides a fully faithful functor
@@
from $\mathbf{AbSch}_{k}$ to the category $\mathbf{Ab}_{k}$ of sheaves of abelian groups on $\mathbf{Aff}_{k}$ endowed with the fppffppf-topology. It is well known that this full embedding commutes with the formations of kernels and cokernels, and that its essential image is also stable by extensions.
@@
The full sub-category of $\mathbf{Ab}_{k}$ formed by sheaves which are representable by commutative group schemes of finite type over $k$, is stable by taking kernels, cokernels and extensions.
Proof. Using [13, Exp. $VI_{A}$ Thm. 5.4.2], and the discussion after [13, Exp. $VI_{A}$ Prop. 5.4.1], we see that the Yoneda embedding is an exact functor. Finally, if a sheaf of abelian groups is an extension of algebraic groups, then it is, by descent, representable by a quasi-separated algebraic space, and we can conclude by [1, Lem. 4.2]. $\Box$
The Barsotti-Chevalley decomposition theorem (see [21]) states that any object $G\in\mathbf{AbSch}_{k}$ fits in an exact sequence
@@
where $G_{0}$ is affine and $A$ is an abelian variety. Moreover, by [12, II-5 Cor. 2.3] $G_{0}$ itself sits in an exact sequence $0\to K\to G_{0}\to H\to 0,$ where $K$ is smooth and affine, and $H$ is finite infinitesimal (i.e. $H_{red}=0$). We also recall from [12, IV-2 Prop. 2.5] that $G\in\mathbf{AbSch}_{k}$ is unipotent if it admits a finite filtration by sub-group schemes
@@
such that each $G_{i}/G_{i+1}$ can be realized as a sub-group scheme of the additive group $\mathbb{G}_{a}$. We can moreover arrange things so that each $G_{i}/G_{i+1}$ is a twisted form of $\mathbb{G}_{a}$, $(\mathbb{Z}/p)^{r}$ for some $r$, or of $\alpha_{p}$. Unipotent group schemes are automatically affine, and form a full sub-category of $\mathbf{AbSch}_{k}$ stable by taking kernels, cokernels and extensions.
@@ 1.2. Reminders on purity for flat cohomology
We remind here the purity theorem we will use in a crucial manner in order to compare algebraic homology with Grothendieck’s original definition. The most general results on purity can be found in [8], but for smooth schemes $X$ over a perfect $k$ these can be proven with simple technology, which we recall below.
@@
Case $G=\mathbb{G}_{m}$. For $d=n$ there is nothing to prove, and when d=n1d=n-1 the statement simply follows from the fact that $X_{x}-\{x\}$ is dense in $X_{x}$. We thus assume that d<n1d<n-1. The case $i=1$ is true by Hartogs, as the restriction map $Pic(X_{x})\to Pic(X_{x}-\{x\})$ is an equivalence of groupoids. For $i=2$ the proposition follows from the fact that $H^{2}_{et}(X_{x},\mathbb{G}_{m})\to H^{2}_{et}(X_{x}-\{x\},\mathbb{G}_{m})$ is always injective (see [15, Cor. 1.10]). When $i>2$, the fact that $\mathbb{G}_{m}$ has torsion cohomology in degree $2$ or higher (see [15, Prop. 1.4]) implies that $H^{i}_{x}(X,\mathbb{G}_{m})$ is always torsion. We are thus led to prove the statement for the group $\mu_{m}$ for some $m$, and by devissage we can even assume that $m$ is prime. If $m$ is invertible in $k$, then the purity theorem in étale cohomology (see for instance [22, §VI Thm. 5.1]) implies the result. When $m=p$ is the characteristic of the base field, we use the following standard "resolution" of $\mu_{p}$: on the small fppf site of $X$, there is an equivalence
@@
where $\Omega_{X/k}^{1,cl}\subset\Omega_{X/k}^{1}$ is the subsheaf of closed $1$-forms and sits in cohomological degrees 0 (see [22, §III Prop. 4.14]). The differential $\Omega_{X/k}^{1,cl}\to\Omega_{X/k}^{1}$ is here iCi-C, where $i$ is the standard inclusion and $C$ is the composition of the projection $\Omega_{X/k}^{1,cl}\to H^{1}_{DR}(X/k)$ and the Cartier isomorphism $C:H^{1}_{DR}(X/k)\simeq\Omega^{1}_{X/k}$. Both abelian sheaves $\Omega_{X/k}^{1,cl}$ and $\Omega_{X/k}^{1}$ are quasi-coherent (the first one on the Frobenius twists $X^{(1)}$ and the second one on $X$), and thus satisfy purity. We deduce from this that $H^{i}_{x}(X,\mu_{p}[1])\simeq H^{i+1}_{x}(X,\mu_{p})\simeq 0$ as soon as i<ndi<n-d and in particular that $H^{i}_{x}(X,\mu_{p})\simeq 0$ when i<ndi<n-d.
@@
1.3. Stacks
We denote by $\mathbf{St}_{k}$ the $\infty$-category of (hypercomplete) stacks over the site $(\mathbf{Aff}_{k},fppf)$, of affine $k$-schemes endowed with the finitely presented flat topology. This $\infty$-category can be described explicitly as being the $\infty$-category of cofibrant and fibrant objects in the model category $\mathbf{Aff}_{k}^{\sim,fppf}$ of presheaves of simplicial sets on $\mathbf{Aff}_{k}$, endowed with the local projective model category structure (see for instance [32, §1.1] and [33, §2.1]). The objects of $\mathbf{Aff}_{k}^{\sim,fppf}$ are functors
$\mathbf{Aff}_{k}^{op}\to sSet,$
from the category of affine $k$-schemes to the category of simplicial sets (we neglect size issues here, see [32] for details concerning universes in this context). For such an object $F$, $X\in\mathbf{Aff}_{k}$, $i>0$ and $s\in F(X)_{0}$ a 0-simplex, we have homotopy sheaves $\pi_{i}(F,s)$, which are sheaves on the site $\mathbf{Aff}_{k}/X$ defined as the fppffppf-sheaf associated to the presheaf $(u:Y\to X)\mapsto\pi_{i}(F(Y),u^{*}(s))$. Similarly, we have a sheaf of connected components $\pi_{0}(F)$, as being associated to the presheaf $X\mapsto\pi_{0}(F(X))$.
The weak equivalences in the model category $\mathbf{Aff}_{k}^{\sim,fppf}$ are the morphisms $F\to G$ inducing isomorphism on all possible homotopy sheaves. The $\infty$-category $\mathbf{St}_{k}$ can then be realized as the $\infty$-category obtained from the genuine category $\mathbf{Aff}_{k}^{\sim,fppf}$ by formally inverting the weak equivalences. A concrete model, up to a natural equivalence of $\infty$-categories, consists of the simplicially enriched category of fibrant and cofibrant objects in $\mathbf{Aff}_{k}^{\sim,fppf}$. Note moreover, that the fibrant objects are precisely the presheaves $F:\mathbf{Aff}_{k}\to sSet$ satisfying the three conditions below.
  1. (1)
    For all $X\in\mathbf{Aff}_{k}$ the simplicial set $F(X)$ satisfies the Kan condition (i.e. is fibrant as a simplicial set).
  2. (2)
    For any finite coproduct of affine schemes $X=\coprod_{i}X_{i}$, the natural morphism $F(X)\to\prod_{i}F(X_{i})$ is an equivalence.
  3. (3)
    For any fppffppf-hypercoverings of affine $k$-schemes $U_{*}\to X$, the natural morphism
    $F(X)\to\underset{[n]\in\Delta}{holim}F(U_{n})$
    is a weak equivalence of simplicial sets.
The last two conditions above form the fppffppf-hyperdescent condition and is the characteristic property of fppffppf-stacks. We refer the reader to [32, 30, 33, 18] for general references on stacks.
We recall the existence of Postnikov decomposition which will be used several times in the sequel. For any stack $F\in\mathbf{St}_{k}$, and any $n\geq 0$, there exists canonical truncations
$p_{n}:F\to\tau_{\leq n}(F)$
which are characterized by the fact that all homotopy sheaves of $\tau_{\leq n}(F)$ vanish in degree strictly higher than $n$, and moreover the morphism $p_{n}$ induces isomorphisms on all homotopy sheaves in degree less or equal to $n$. For any $n>1$, there exists a cartesian square in $\mathbf{St}_{k}/\tau_{\leq 1}(F)$
${\tau_{\leq n}(F)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\tau_{\leq n-1}(F)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k_{n}}\textstyle{\bullet\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K(\pi_{n}(F),n+1)}$
where $\pi_{i}(F)$ is the $n$-th homotopy sheaf of $F$, which is a sheaf of abelian groups over the $1$-truncation $\tau_{\leq 1}(F)$. Here $K(\pi_{n}(F),n+1)$ denotes the Eilenberg-MacLane stack (relative to $\tau_{\leq 1}(F)$) whose unique non-zero homotopy sheaf is $\pi_{n}(F)$ in degree $n+1$. The morphism $k_{n}$ is called the $n$-th Postnikov invariant of $F$. In our context, the stacks considered will always be abelian group stacks, and thus the sheaves $\pi_{n}(F)$ will be constant over $\tau_{\leq 1}(F)$ and simply be considered as fppffppf-sheaves over $\mathbf{Aff}_{k}$.